Suppose that the following summation notation represents a geometric series: sum j=1 ^ n a If the ratio of this series is r, which of the following represent the sum of the series? Check all that apply
1
Expert's answer
2019-10-24T07:02:58-0400

"\sum\limits_{j=1}^n a_j"

Since

"a_j=a_{j-1}r=a_1r^{j-1}, r\ne 1"

we assume that

"\sum\limits_{j=1}^n a_j=\sum\limits_{j=1}^n a_1r^{j-1}=a_1\sum\limits_{j=1}^n r^{j-1}"

It is known, that

"(1-r^n)=(1-r)(1+r+r^2+...+r^{n-1})"

Hence

"\sum\limits_{j=1}^n r^{j-1}=\frac{1-r^n}{1-r}"

Therefore

"\sum\limits_{j=1}^n a_j=a_1\frac{1-r^n}{1-r}"

To find the sum of the series we need to find

"\lim\limits_{n\to \infty}\sum\limits_{j=1}^n a_j=\lim\limits_{n\to \infty}a_1\frac{1-r^n}{1-r}=\left\{
\begin{matrix}
 \infty , & r>1 \\
 \frac{a_1}{1-r}, & r<1
\end{matrix} 
\right."